# Efficient designs for multivariate crossover trials

### v1

@@ A.2  Generalized markov-type structure
Here $d^{*}\in\mathcal{D}^{(1)}_{t,n=\lambda t(t-1),p=t}$ be a design given by $OA_{I}\big{(}n=\lambda t\left(t-1\right),p=t,$
$t,2\big{)}$, where $\lambda$ is a positive integer and $t\geq 3$, and thus $d^{*}$ is a design uniform on periods. Since from Remark S.1, we get that $\bm{{1}}_{p}^{{}^{\prime}}\bm{\varOmega}_{1}=\bm{{1}}_{p}^{{}^{\prime}}\bm{\varOmega}_{2}=\bm{{1}}_{p}^{{}^{\prime}}\bm{\varOmega}_{4}=\bm{{0}}_{1\times p}$, $\bm{{1}}_{p}^{{}^{\prime}}\bm{\psi}^{{}^{\prime}}\bm{\varOmega}_{x}\bm{\psi}\bm{{1}}_{p}$ is the entry corresponding to the first row and first column of the matrix $\bm{\varOmega}_{x}$, for $x=1,2,4$. Here, $\bm{\varOmega}_{1}$, $\bm{\varOmega}_{2}$ and $\bm{\varOmega}_{4}$ are as given in Lemma S.4. Using the expression of $\bm{{C}}_{d^{*}(s2)}$ as given in Lemma S.3 and the expression of $\bm{A}^{*}$ in Lemma S.4, we get
@@
$\begin{split}\bm{C}_{{d^{*}(s2)}(11)}&=\begin{bmatrix}\bm{T}_{d^{*}}^{{}^{\prime}}\left(\bm{H}_{n}\otimes\bm{\varOmega}_{1}\right)\bm{T}_{d^{*}}&-\bm{T}_{d^{*}}^{{}^{\prime}}\left(\bm{H}_{n}\otimes\bm{\varOmega}_{2}\right)\bm{T}_{d^{*}}\\ -\bm{T}_{d^{*}}^{{}^{\prime}}\left(\bm{H}_{n}\otimes\bm{\varOmega}_{2}\right)\bm{T}_{d^{*}}&\bm{T}_{d^{*}}^{{}^{\prime}}\left(\bm{H}_{n}\otimes\bm{\varOmega}_{4}\right)\bm{T}_{d^{*}}\end{bmatrix},\\ \bm{C}_{{d^{*}(s2)}(12)}&=\bm{C}^{{}^{\prime}}_{{d^{*}(s2)}(21)}=\begin{bmatrix}\bm{T}_{d^{*}}^{{}^{\prime}}\left(\bm{H}_{n}\otimes\bm{\varOmega}_{1}\bm{\psi}\right)\bm{T}_{d^{*}}&-\bm{T}_{d^{*}}^{{}^{\prime}}\left(\bm{H}_{n}\otimes\bm{\varOmega}_{2}\bm{\psi}\right)\bm{T}_{d^{*}}\\ -\bm{T}_{d^{*}}^{{}^{\prime}}\left(\bm{H}_{n}\otimes\bm{\varOmega}_{2}\bm{\psi}\right)\bm{T}_{d^{*}}&\bm{T}_{d^{*}}^{{}^{\prime}}\left(\bm{H}_{n}\otimes\bm{\varOmega}_{4}\bm{\psi}\right)\bm{T}_{d^{*}}\end{bmatrix},\text{ and}\\ \bm{C}_{{d^{*}(s2)}(22)}&=\begin{bmatrix}\bm{T}_{d^{*}}^{{}^{\prime}}\left(\bm{H}_{n}\otimes\bm{\psi}^{{}^{\prime}}\bm{\varOmega}_{1}\bm{\psi}\right)\bm{T}_{d^{*}}&-\bm{T}_{d^{*}}^{{}^{\prime}}\left(\bm{H}_{n}\otimes\bm{\psi}^{{}^{\prime}}\bm{\varOmega}_{2}\bm{\psi}\right)\bm{T}_{d^{*}}\\ -\bm{T}_{d^{*}}^{{}^{\prime}}\left(\bm{H}_{n}\otimes\bm{\psi}^{{}^{\prime}}\bm{\varOmega}_{2}\bm{\psi}\right)\bm{T}_{d^{*}}&\bm{T}_{d^{*}}^{{}^{\prime}}\left(\bm{H}_{n}\otimes\bm{\psi}^{{}^{\prime}}\bm{\varOmega}_{4}\bm{\psi}\right)\bm{T}_{d^{*}}\end{bmatrix}.\end{split}$
@@
where $c_{11(11)}=tr\left(\bm{\varOmega}_{1}\right)$, $c_{11(12)}=tr\left(\bm{\varOmega}_{2}\right)$, $c_{11(22)}=tr\left(\bm{\varOmega}_{4}\right)$, $c_{12(11)}=tr\left(\bm{\varOmega}_{1}\bm{\psi}\right)$, $c_{12(12)}=tr\left(\bm{\varOmega}_{2}\bm{\psi}\right)$, $c_{12(22)}=tr\left(\bm{\varOmega}_{4}\bm{\psi}\right)$, $c_{22(11)}=tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{\varOmega}_{1}\bm{\psi}\right)$, $c_{22(12)}=tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{\varOmega}_{2}\bm{\psi}\right)$, $c_{22(22)}=tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{\varOmega}_{4}\bm{\psi}\right)$. From Lemma S.4, we get
$\begin{split}\bm{\varOmega}_{1}&=\bm{V}_{1}^{*}+\frac{\bar{\rho}^{2}}{\sigma_{12}}\bm{V}_{R}^{*}=\bm{V}_{1}^{*}+\bar{\rho}^{2}\bm{\varOmega}_{4},\\ \bm{\varOmega}_{2}&=\bar{\rho}\bm{\varOmega}_{4},\text{ and}\\ \bm{\varOmega}_{4}&=\frac{1}{\sigma_{12}}\bm{V}_{R}^{*},\end{split}$ (A.8)
@@
$tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{V}_{R}^{*}\bm{\psi}\right)$. Since $\sigma_{12}>0$, applying Lemma S.8, we get that $c_{22(11)}c_{22(22)}-c^{2}_{22(12)}>0$. So, from (A.7), we get
@@
From (A.6), (A.7), (A.9) and (A.12), by calculation we get that $\bm{{C}}_{d^{*}(s2)}=$
$\frac{n}{t-1}\begin{bmatrix}\bm{\varLambda}_{1}&\bm{\varLambda}_{2}\\ \bm{\varLambda}_{2}&\bm{\varLambda}_{4}\end{bmatrix}$, where
$\begin{split}\bm{\varLambda}_{1}&=\left(tr\left(\bm{V}_{1}^{*}\right)+\frac{\bar{\rho}^{2}}{\sigma_{12}}tr\left(\bm{V}_{R}^{*}\right)-\frac{\left(tr\left(\bm{V}_{1}^{*}\bm{\psi}\right)\right)^{2}}{tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{V}_{1}^{*}\bm{\psi}\right)}-\frac{\bar{\rho}^{2}}{\sigma_{12}}\frac{\left(tr\left(\bm{V}_{R}^{*}\bm{\psi}\right)\right)^{2}}{tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{V}_{R}^{*}\bm{\psi}\right)}\right)\bm{H}_{t},\\ \bm{\varLambda}_{2}&=-\frac{\bar{\rho}}{\sigma_{12}}\left(tr\left(\bm{V}_{R}^{*}\right)-\frac{\left(tr\left(\bm{V}_{R}^{*}\bm{\psi}\right)\right)^{2}}{tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{V}_{R}^{*}\bm{\psi}\right)}\right)\bm{H}_{t},\text{ and}\\ \bm{\varLambda}_{4}&=\frac{1}{\sigma_{12}}\left(tr\left(\bm{V}_{R}^{*}\right)-\frac{\left(tr\left(\bm{V}_{R}^{*}\bm{\psi}\right)\right)^{2}}{tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{V}_{R}^{*}\bm{\psi}\right)}\right)\bm{H}_{t}.\end{split}$
@@
$\bm{{C}}_{d^{*}(s2)}$ $=\frac{n}{t-1}\begin{bmatrix}\bm{\varLambda}_{1}&\bm{\varLambda}_{2}\\ \bm{\varLambda}_{2}&\bm{\varLambda}_{4}\end{bmatrix},$ (A.15)
@@
$\begin{split}\bm{\varLambda}_{1}&=\left(tr\left(\bm{V}_{1}^{*}\right)+\frac{\bar{\rho}^{2}}{\sigma_{12}}tr\left(\bm{V}_{R}^{*}\right)-\frac{\left(tr\left(\bm{V}_{1}^{*}\bm{\psi}\right)\right)^{2}}{tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{V}_{1}^{*}\bm{\psi}\right)}-\frac{\bar{\rho}^{2}}{\sigma_{12}}\frac{\left(tr\left(\bm{V}_{R}^{*}\bm{\psi}\right)\right)^{2}}{tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{V}_{R}^{*}\bm{\psi}\right)}\right)\bm{H}_{t},\\ \bm{\varLambda}_{2}&=-\frac{\bar{\rho}}{\sigma_{12}}\left(tr\left(\bm{V}_{R}^{*}\right)-\frac{\left(tr\left(\bm{V}_{R}^{*}\bm{\psi}\right)\right)^{2}}{tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{V}_{R}^{*}\bm{\psi}\right)}\right)\bm{H}_{t},\text{ and}\\ \bm{\varLambda}_{4}&=\frac{1}{\sigma_{12}}\left(tr\left(\bm{V}_{R}^{*}\right)-\frac{\left(tr\left(\bm{V}_{R}^{*}\bm{\psi}\right)\right)^{2}}{tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{V}_{R}^{*}\bm{\psi}\right)}\right)\bm{H}_{t}.\end{split}$
@@ Appendix B Efficiency under the generalized markov-type structure
Here, $d^{*}\in\mathcal{D}^{(1)}_{t,n=\lambda t(t-1),p=t}$ is a design given by $OA_{I}\big{(}n=\lambda t\left(t-1\right),p=t,$
$t,2\big{)}$, where $\lambda$ is a positive integer and $t\geq 3$. Using the expression of $tr\left(\bm{{C}}_{d^{*}(s2)}\right)$ as given in Theorem 3.3 and the expression of $u\left(t,n,p,\sigma_{11},\sigma_{22},\rho,\bm{V}_{1},\bm{V}_{R}\right)$ as given in Lemma 3.1, where $p=t$ and $n=\lambda t\left(t-1\right)$, by calculation we get
@@
where $\bm{V}_{C}^{*}=\bm{V}_{C}^{-1}-\left(\bm{{1}}_{p}^{{}^{\prime}}\bm{V}_{C}^{-1}\bm{{1}}_{p}\right)^{-1}\bm{V}_{C}^{-1}\bm{J}_{p\times p}\bm{V}_{C}^{-1}$ and $\bm{V}_{R}^{*}=\bm{V}_{R}^{-1}-\left(\bm{{1}}_{p}^{{}^{\prime}}\bm{V}_{R}^{-1}\bm{{1}}_{p}\right)^{-1}\bm{V}_{R}^{-1}\bm{J}_{p\times p}\bm{V}_{R}^{-1}$, and using the expressions of $\bm{\varOmega}_{1}$, $\bm{\varOmega}_{2}$ and $\bm{\varOmega}_{4}$ as given in Lemma S.4, we get
$\begin{split}\bm{\varOmega}_{1}&=\frac{1}{\sigma_{11}}\bm{V}^{*}_{C}+\frac{\rho^{2}}{\sigma_{11}\left(1-\rho^{2}\right)}\bm{V}_{R}^{*},\\ \bm{\varOmega}_{2}&=\frac{\rho}{\sigma_{11}\left(1-\rho^{2}\right)}\bm{V}_{R}^{*}\\ \bm{\varOmega}_{4}&=\frac{1}{\sigma_{11}\left(1-\rho^{2}\right)}\bm{V}_{R}^{*}\end{split}$ (B.3)
@@
Gene Marker
Characteristics
PNO1 (X100134584_TGI_AT)
Facilitates ribosomal RNA maturation and processing, pivotal for ribosome biogenesis and protein synthesis control.
BPI (X100126967_TGI_AT)
Defends against bacterial infections by neutralizing bacterial endotoxins within the innate immune system.
UNC13A (X100308708_TGI_AT)
Regulates neurotransmitter release, contributing to synaptic plasticity and efficient neuronal communication.
MTA3 (X100154937_TGI_AT)
Exhibits chromatin-binding and histone deacetylase activities, influencing transcriptional regulation and chromatin remodeling.
DYRK1A (X100140918_TGI_AT)
Multifaceted kinase with roles in circadian rhythm, neuronal development, and protein phosphorylation-mediated cellular processes.
EPN3 (X100310234_TGI_AT)
Involved in lipid and protein binding, localized in vital cellular compartments, including clathrin-coated pits and plasma membranes.
ZNF365 (X100300789_TGI_AT)
Participates in DNA binding and transcriptional regulation, impacting gene expression and diverse biological contexts.
OR5AS1 (X100125346_TGI_AT)
Essential component of the olfactory receptor family, contributing to chemosensory perception and the sense of smell.
EHMT1 (X100129988_TGI_AT)
Epigenetic regulator governing histone methylation, chromatin modification, and gene expression, with implications in embryonic development and disorders.
LTBP2 (X100129645_TGI_AT)
Crucial in the extracellular matrix, regulating transforming growth factor-beta (TGF-$\beta$) activity, influencing tissue homeostasis, and cellular behavior.
Table S.1: Characteristics of the $10$ gene markers. This information is obtained from GEO’s annotation SOFT table.
@@
where $\bm{C}_{d(s1)(11)}=\left(\bm{I}_{g}\otimes\bm{T}^{{}^{\prime}}_{d}\right)\bm{A}^{*}\left(\bm{I}_{g}\otimes\bm{T}_{d}\right)$, $\bm{C}_{d(s1)(12)}=\bm{C}^{{}^{\prime}}_{d(s1)(21)}=\left(\bm{I}_{g}\otimes\bm{T}^{{}^{\prime}}_{d}\right)\bm{A}^{*}\left(\bm{I}_{g}\otimes\bm{F}_{d}\right)$ and $\bm{C}_{d(s1)(22)}=\left(\bm{I}_{g}\otimes\bm{F}^{{}^{\prime}}_{d}\right)\bm{A}^{*}\left(\bm{I}_{g}\otimes\bm{F}_{d}\right)$. Here, $\bm{A}^{*}=\bm{\varSigma}^{-1/2}pr^{\perp}\left(\bm{\varSigma}^{-1/2}\bm{Z}_{1}\right)\bm{\varSigma}^{-1/2}$.

### v2

@@ A.2  Generalized Markov-type structure
Here $d^{*}\in\mathcal{D}^{(1)}_{t,n=\lambda t(t-1),p=t}$ be a design given by $OA_{I}\big{(}n=\lambda t\left(t-1\right),p=t,t,2\big{)}$, where $\lambda$ is a positive integer and $t\geq 3$, and thus $d^{*}$ is a design uniform on periods. Since from Remark S.1, we get that $\bm{{1}}_{p}^{{}^{\prime}}\bm{\mathit{\Omega}}_{1}=\bm{{1}}_{p}^{{}^{\prime}}\bm{\mathit{\Omega}}_{2}=\bm{{1}}_{p}^{{}^{\prime}}\bm{\mathit{\Omega}}_{4}=\bm{{0}}_{1\times p}$, $\bm{{1}}_{p}^{{}^{\prime}}\bm{\psi}^{{}^{\prime}}\bm{\mathit{\Omega}}_{x}\bm{\psi}\bm{{1}}_{p}$ is the entry corresponding to the first row and first column of the matrix $\bm{\mathit{\Omega}}_{x}$, for $x=1,2,4$. Here $\bm{\mathit{\Omega}}_{1}$, $\bm{\mathit{\Omega}}_{2}$ and $\bm{\mathit{\Omega}}_{4}$ are as given in Lemma S.4. Using the expression of $\bm{{C}}_{d^{*}(s2)}$ as given in Lemma S.3 and the expression of $\bm{A}^{*}$ in Lemma S.4, we get
@@
$\begin{split}\bm{C}_{{d^{*}(s2)}(11)}&=\begin{bmatrix}\bm{T}_{d^{*}}^{{}^{\prime}}\left(\bm{H}_{n}\otimes\bm{\mathit{\Omega}}_{1}\right)\bm{T}_{d^{*}}&-\bm{T}_{d^{*}}^{{}^{\prime}}\left(\bm{H}_{n}\otimes\bm{\mathit{\Omega}}_{2}\right)\bm{T}_{d^{*}}\\ -\bm{T}_{d^{*}}^{{}^{\prime}}\left(\bm{H}_{n}\otimes\bm{\mathit{\Omega}}_{2}\right)\bm{T}_{d^{*}}&\bm{T}_{d^{*}}^{{}^{\prime}}\left(\bm{H}_{n}\otimes\bm{\mathit{\Omega}}_{4}\right)\bm{T}_{d^{*}}\end{bmatrix},\\ \bm{C}_{{d^{*}(s2)}(12)}&=\bm{C}^{{}^{\prime}}_{{d^{*}(s2)}(21)}=\begin{bmatrix}\bm{T}_{d^{*}}^{{}^{\prime}}\left(\bm{H}_{n}\otimes\bm{\mathit{\Omega}}_{1}\bm{\psi}\right)\bm{T}_{d^{*}}&-\bm{T}_{d^{*}}^{{}^{\prime}}\left(\bm{H}_{n}\otimes\bm{\mathit{\Omega}}_{2}\bm{\psi}\right)\bm{T}_{d^{*}}\\ -\bm{T}_{d^{*}}^{{}^{\prime}}\left(\bm{H}_{n}\otimes\bm{\mathit{\Omega}}_{2}\bm{\psi}\right)\bm{T}_{d^{*}}&\bm{T}_{d^{*}}^{{}^{\prime}}\left(\bm{H}_{n}\otimes\bm{\mathit{\Omega}}_{4}\bm{\psi}\right)\bm{T}_{d^{*}}\end{bmatrix},\text{ and}\\ \bm{C}_{{d^{*}(s2)}(22)}&=\begin{bmatrix}\bm{T}_{d^{*}}^{{}^{\prime}}\left(\bm{H}_{n}\otimes\bm{\psi}^{{}^{\prime}}\bm{\mathit{\Omega}}_{1}\bm{\psi}\right)\bm{T}_{d^{*}}&-\bm{T}_{d^{*}}^{{}^{\prime}}\left(\bm{H}_{n}\otimes\bm{\psi}^{{}^{\prime}}\bm{\mathit{\Omega}}_{2}\bm{\psi}\right)\bm{T}_{d^{*}}\\ -\bm{T}_{d^{*}}^{{}^{\prime}}\left(\bm{H}_{n}\otimes\bm{\psi}^{{}^{\prime}}\bm{\mathit{\Omega}}_{2}\bm{\psi}\right)\bm{T}_{d^{*}}&\bm{T}_{d^{*}}^{{}^{\prime}}\left(\bm{H}_{n}\otimes\bm{\psi}^{{}^{\prime}}\bm{\mathit{\Omega}}_{4}\bm{\psi}\right)\bm{T}_{d^{*}}\end{bmatrix}.\end{split}$
@@
where $c_{11(11)}=tr\left(\bm{\mathit{\Omega}}_{1}\right)$, $c_{11(12)}=tr\left(\bm{\mathit{\Omega}}_{2}\right)$, $c_{11(22)}=tr\left(\bm{\mathit{\Omega}}_{4}\right)$, $c_{12(11)}=tr\left(\bm{\mathit{\Omega}}_{1}\bm{\psi}\right)$, $c_{12(12)}=tr\left(\bm{\mathit{\Omega}}_{2}\bm{\psi}\right)$, $c_{12(22)}=tr\left(\bm{\mathit{\Omega}}_{4}\bm{\psi}\right)$, $c_{22(11)}=tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{\mathit{\Omega}}_{1}\bm{\psi}\right)$, $c_{22(12)}=tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{\mathit{\Omega}}_{2}\bm{\psi}\right)$, $c_{22(22)}=tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{\mathit{\Omega}}_{4}\bm{\psi}\right)$. From Lemma S.4, we get
$\begin{split}\bm{\mathit{\Omega}}_{1}&=\bm{V}_{1}^{*}+\frac{\bar{\rho}^{2}}{\sigma_{12}}\bm{V}_{R}^{*}=\bm{V}_{1}^{*}+\bar{\rho}^{2}\bm{\mathit{\Omega}}_{4},\\ \bm{\mathit{\Omega}}_{2}&=\bar{\rho}\bm{\mathit{\Omega}}_{4},\text{ and}\\ \bm{\mathit{\Omega}}_{4}&=\frac{1}{\sigma_{12}}\bm{V}_{R}^{*},\end{split}$ (A.8)
@@
$tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{V}_{R}^{*}\bm{\psi}\right)$. Since $\sigma_{12}>0$, applying Lemma S.8, we get that $c_{22(11)}c_{22(22)}-c^{2}_{22(12)}>0$. So from (A.7), we get
@@
From (A.6), (A.7), (A.9) and (A.10), by calculation we get that $\bm{{C}}_{d^{*}(s2)}=\frac{n}{t-1}\begin{bmatrix}\bm{\mathit{\Lambda}}_{1}&\bm{\mathit{\Lambda}}_{2}\\ \bm{\mathit{\Lambda}}_{2}&\bm{\mathit{\Lambda}}_{4}\end{bmatrix}$, where
$\begin{split}\bm{\mathit{\Lambda}}_{1}&=\left(tr\left(\bm{V}_{1}^{*}\right)+\frac{\bar{\rho}^{2}}{\sigma_{12}}tr\left(\bm{V}_{R}^{*}\right)-\frac{\left(tr\left(\bm{V}_{1}^{*}\bm{\psi}\right)\right)^{2}}{tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{V}_{1}^{*}\bm{\psi}\right)}-\frac{\bar{\rho}^{2}}{\sigma_{12}}\frac{\left(tr\left(\bm{V}_{R}^{*}\bm{\psi}\right)\right)^{2}}{tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{V}_{R}^{*}\bm{\psi}\right)}\right)\bm{H}_{t},\\ \bm{\mathit{\Lambda}}_{2}&=-\frac{\bar{\rho}}{\sigma_{12}}\left(tr\left(\bm{V}_{R}^{*}\right)-\frac{\left(tr\left(\bm{V}_{R}^{*}\bm{\psi}\right)\right)^{2}}{tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{V}_{R}^{*}\bm{\psi}\right)}\right)\bm{H}_{t},\text{ and}\\ \bm{\mathit{\Lambda}}_{4}&=\frac{1}{\sigma_{12}}\left(tr\left(\bm{V}_{R}^{*}\right)-\frac{\left(tr\left(\bm{V}_{R}^{*}\bm{\psi}\right)\right)^{2}}{tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{V}_{R}^{*}\bm{\psi}\right)}\right)\bm{H}_{t}.\end{split}$
@@
$\bm{{C}}_{d^{*}(s2)}$ $=\frac{n}{t-1}\begin{bmatrix}\bm{\mathit{\Lambda}}_{1}&\bm{\mathit{\Lambda}}_{2}\\ \bm{\mathit{\Lambda}}_{2}&\bm{\mathit{\Lambda}}_{4}\end{bmatrix},$ (A.11)
@@
$\begin{split}\bm{\mathit{\Lambda}}_{1}&=\left(tr\left(\bm{V}_{1}^{*}\right)+\frac{\bar{\rho}^{2}}{\sigma_{12}}tr\left(\bm{V}_{R}^{*}\right)-\frac{\left(tr\left(\bm{V}_{1}^{*}\bm{\psi}\right)\right)^{2}}{tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{V}_{1}^{*}\bm{\psi}\right)}-\frac{\bar{\rho}^{2}}{\sigma_{12}}\frac{\left(tr\left(\bm{V}_{R}^{*}\bm{\psi}\right)\right)^{2}}{tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{V}_{R}^{*}\bm{\psi}\right)}\right)\bm{H}_{t},\\ \bm{\mathit{\Lambda}}_{2}&=-\frac{\bar{\rho}}{\sigma_{12}}\left(tr\left(\bm{V}_{R}^{*}\right)-\frac{\left(tr\left(\bm{V}_{R}^{*}\bm{\psi}\right)\right)^{2}}{tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{V}_{R}^{*}\bm{\psi}\right)}\right)\bm{H}_{t},\text{ and}\\ \bm{\mathit{\Lambda}}_{4}&=\frac{1}{\sigma_{12}}\left(tr\left(\bm{V}_{R}^{*}\right)-\frac{\left(tr\left(\bm{V}_{R}^{*}\bm{\psi}\right)\right)^{2}}{tr\left(\bm{H}_{p}\bm{\psi}^{{}^{\prime}}\bm{V}_{R}^{*}\bm{\psi}\right)}\right)\bm{H}_{t}.\end{split}$
@@ Appendix B Efficiency under the generalized Markov-type structure
Here, $d^{*}\in\mathcal{D}^{(1)}_{t,n=\lambda t(t-1),p=t}$ is a design given by $OA_{I}\big{(}n=\lambda t\left(t-1\right),p=t,t,2\big{)}$, where $\lambda$ is a positive integer and $t\geq 3$. Using the expression of $tr\left(\bm{{C}}_{d^{*}(s2)}\right)$ as given in Theorem 3.3 and the expression of $u\left(t,n,p,\sigma_{11},\sigma_{22},\rho,\bm{V}_{1},\bm{V}_{R}\right)$ as given in Lemma 3.1, where $p=t$ and $n=\lambda t\left(t-1\right)$, by calculation we get
@@
where $\bm{V}_{C}^{*}=\bm{V}_{C}^{-1}-\left(\bm{{1}}_{p}^{{}^{\prime}}\bm{V}_{C}^{-1}\bm{{1}}_{p}\right)^{-1}\bm{V}_{C}^{-1}\bm{J}_{p\times p}\bm{V}_{C}^{-1}$ and $\bm{V}_{R}^{*}=\bm{V}_{R}^{-1}-\left(\bm{{1}}_{p}^{{}^{\prime}}\bm{V}_{R}^{-1}\bm{{1}}_{p}\right)^{-1}\bm{V}_{R}^{-1}\bm{J}_{p\times p}\bm{V}_{R}^{-1}$, and using the expressions of $\bm{\mathit{\Omega}}_{1}$, $\bm{\mathit{\Omega}}_{2}$ and $\bm{\mathit{\Omega}}_{4}$ as given in Lemma S.4, we get
$\begin{split}\bm{\mathit{\Omega}}_{1}&=\frac{1}{\sigma_{11}}\bm{V}^{*}_{C}+\frac{\rho^{2}}{\sigma_{11}\left(1-\rho^{2}\right)}\bm{V}_{R}^{*},\\ \bm{\mathit{\Omega}}_{2}&=\frac{\rho}{\sigma_{11}\left(1-\rho^{2}\right)}\bm{V}_{R}^{*},\text{ and}\\ \bm{\mathit{\Omega}}_{4}&=\frac{1}{\sigma_{11}\left(1-\rho^{2}\right)}\bm{V}_{R}^{*}.\end{split}$ (B.3)
@@
Declaration of competing interest
The authors declare no conflicts of interest.
Funding
This research did not receive any specific grant from funding agencies in the public, commercial, or not-for-profit sectors.
Author contributions
Shubham Niphadkar: Formal analysis, Methodology, Software, Writing-Original draft preparation and Editing. Siuli Mukhopadhyay: Conceptualization, Methodology, Validation, Writing-Reviewing and Editing.
@@
       Gene Marker        Characteristics
       BPI (X100126967_TGI_AT)        Defends against bacterial infections by neutralizing bacterial endotoxins within the innate immune system.
       MTA3 (X100154937_TGI_AT)        Exhibits chromatin-binding and histone deacetylase activities, influencing transcriptional regulation and chromatin remodeling.
       LTBP2 (X100129645_TGI_AT)        Crucial in the extracellular matrix, regulating transforming growth factor-beta (TGF-$\beta$) activity, influencing tissue homeostasis, and cellular behavior.
Table S.1: Characteristics of the $3$ gene markers. This information is obtained from GEO’s annotation SOFT table.
@@
where $\bm{C}_{d(s1)(11)}=\left(\bm{I}_{g}\otimes\bm{T}^{{}^{\prime}}_{d}\right)\bm{A}^{*}\left(\bm{I}_{g}\otimes\bm{T}_{d}\right)$, $\bm{C}_{d(s1)(12)}=\bm{C}^{{}^{\prime}}_{d(s1)(21)}=\left(\bm{I}_{g}\otimes\bm{T}^{{}^{\prime}}_{d}\right)\bm{A}^{*}\left(\bm{I}_{g}\otimes\bm{F}_{d}\right)$ and $\bm{C}_{d(s1)(22)}=\left(\bm{I}_{g}\otimes\bm{F}^{{}^{\prime}}_{d}\right)\bm{A}^{*}\left(\bm{I}_{g}\otimes\bm{F}_{d}\right)$. Here, $\bm{A}^{*}=\bm{\mathit{\Sigma}}^{-1/2}pr^{\perp}\left(\bm{\mathit{\Sigma}}^{-1/2}\bm{Z_{1}}\right)\bm{\mathit{\Sigma}}^{-1/2}$.